# The numeraire e-variable

### 29 Feb 2024 v1

@@ 4 Finding the numeraire
We identify ${\mathcal{P}}$ with its set of densities $C=\{p\in L^{0}_{+}(\mu)\colon p=d{\mathbb{P}}/d\mu,\ {\mathbb{P}}\in{\mathcal{P}}\}$. The polar of $C$ in $L^{0}_{+}(\mu)$ is $C^{\circ}=\{f\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }p\in C\}$, which is the set of (equivalence classes of) random variables that are $\mu$-almost surely equal to a finite e-variable. We claim that all elements of ${\mathcal{P}}^{\circ\circ}$ are absolutely continuous with respect to $\mu$, and that its set of densities is precisely the bipolar $C^{\circ\circ}=\{p\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }f\in C^{\circ}\}$. To check absolute continuity, let ${\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$ and consider any event $A$ such that $\mu(A)=0$. Then $\infty\bm{1}_{A}$ is an e-variable, hence ${\mathbb{E}}_{\mathbb{P}}[\infty\bm{1}_{A}]\leq 1$, and thus ${\mathbb{P}}(A)=0$. This shows absolute continuity and lets us identify ${\mathcal{P}}^{\circ\circ}$ with its set of densities, namely, the set of $p\in L^{0}_{+}(\mu)$ such that ${\mathbb{E}}_{\mu}[pX]\leq 1$ for all finite e-variables $X$. Here we may restrict to finite e-variables $X$ because ${\mathbb{E}}_{\mu}[pX]\leq 1$ if and only if ${\mathbb{E}}_{\mu}[p(X\wedge n)]\leq 1$ for all $n\in{\mathbb{N}}$, and each $X\wedge n$ is a finite e-variable. These are, up to $\mu$-nullsets, precisely the elements of $C^{\circ}$, so we deduce that ${\mathcal{P}}^{\circ\circ}$ is indeed identified with $C^{\circ\circ}$. Finally, the Bipolar Theorem 1.3 in Brannath and Schachermayer (1999) states that $C^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains $C$. This completes the proof. ∎
@@
If a reference measure $\mu$ is given, we identify any ${\mathbb{P}}\in{\mathcal{P}}$ with its density $p=d{\mathbb{P}}/d\mu$. Thanks to Lemma 4.5, we may similarly identify ${\mathcal{P}}^{\circ\circ}$ with its set of densities. For the sake of brevity, we abuse notation and write, for example, $p\in{\mathcal{P}}$ even though, strictly speaking, ${\mathcal{P}}$ is a set of measures, not densities. Next, given ${\mathbb{Q}}$ with density $q$, the absolutely continuous part of any $p\in{\mathcal{P}}^{\circ\circ}$ has density $p^{a}=p\bm{1}_{\{q>0\}}$, and this is ${\mathbb{Q}}$-almost surely equal to $p$ itself. In view of these observations, the conditions (3.3) and (3.4) of Theorem 3.4 characterizing the RIPr state that ${\mathbb{E}}_{\mathbb{Q}}[p/p^{*}]\leq 1$, respectively ${\mathbb{E}}_{\mathbb{Q}}[\log(p/p^{*})]\leq 0$, for all $p\in{\mathcal{P}}^{\circ\circ}$. Here $p^{*}=d{\mathbb{P}}^{*}/d\mu$ is the density of the RIPr. The following theorem simplifies these conditions by showing that, essentially, it is enough to check them for elements of ${\mathcal{P}}$ rather than ${\mathcal{P}}^{\circ\circ}$.
@@
Assume that ${\mathbb{Q}}\ll{\mathcal{P}}$ and that ${\mathbb{Q}}$ and all elements of ${\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Let $p^{*}\in{\mathcal{P}}^{\circ\circ}$ be equivalent to ${\mathbb{Q}}$. Then each of the conditions (3.3) and (3.4) is equivalent to
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{p}{p^{*}}\right]\leq 1\text{ for all $p\in{\mathcal{P}}$,}$ (4.3)
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\log\frac{(1-t)p^{*}+tp}{p^{*}}\right]\leq 0\text{ for all $t\in(0,1)$ and $p\in{\mathcal{P}}$.}$ (4.4)
@@
The reason for the asymmetry between (3.4) and (4.4) is that the latter only involves $p$ from ${\mathcal{P}}$, not ${\mathcal{P}}^{\circ\circ}$. Since ${\mathcal{P}}$ may not contain $p^{*}$, nor be convex, we need to explicitly include convex combinations with $p^{*}$ in the numerator on the left-hand side of (4.4). Let us also emphasize that $t=1$ is excluded in (4.4). This sometimes makes the condition easier to check, for instance in the proof of Corollary 4.8 below.
@@
We know from Theorem 3.4 that (3.3) and (3.4) are equivalent and uniquely characterize the RIPr among the elements of ${\mathcal{P}}^{\circ\circ}$ that are equivalent to ${\mathbb{Q}}$. To prove the theorem it is therefore enough to show that that (3.3), (4.3), and (4.4) are equivalent.
We first show that (3.3) is equivalent to (4.3). The former is clearly a stronger condition because it states that (4.3) holds with the larger set ${\mathcal{P}}^{\circ\circ}$ in place of ${\mathcal{P}}$. For the converse, suppose (4.3) holds. Let ${\mathcal{P}}^{\prime}$ be the set of all sub-probability densities dominated by some element of the convex hull of ${\mathcal{P}}$, and let ${\mathcal{P}}^{\prime\prime}$ be the closure in $\mu$-probability of ${\mathcal{P}}^{\prime}$. It is clear that (4.3) holds for all $p\in{\mathcal{P}}^{\prime}$ and then, by Fatou’s lemma, also for all $p\in{\mathcal{P}}^{\prime\prime}$. We claim that ${\mathcal{P}}^{\prime\prime}$ is a closed convex solid set that contains ${\mathcal{P}}$; it must then also contain ${\mathcal{P}}^{\circ\circ}$ since this is the smallest such set by Lemma 4.5. From this we conclude that (3.3) holds. To see that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid, note that by definition, every $p^{\prime\prime}\in{\mathcal{P}}^{\prime\prime}$ is the limit in probability of some $p_{n}^{\prime}\in{\mathcal{P}}^{\prime}$, meaning that $p_{n}^{\prime}\leq p_{n}$ for some $p_{n}$ in the convex hull of ${\mathcal{P}}$. From this description convexity follows directly, closedness is obtained from a diagonal argument, and solidity follows by noting that any $\widetilde{p}\leq p^{\prime\prime}$ is the limit of $\widetilde{p}\wedge p_{n}^{\prime}$, which is still dominated by $p_{n}$ and thus belongs to ${\mathcal{P}}^{\prime}$.
The equivalence of (4.3) and (4.4) is argued as the equivalence of (3.3) and (3.4), or the equivalence of the numeraire property and log-optimality in Proposition 2.4. Specifically, the forward implication uses Jensen’s inequality and (4.3) to get ${\mathbb{E}}_{\mathbb{Q}}[\log(1-t+tp/p^{*})]\leq\log(1-t+t{\mathbb{E}}_{\mathbb{Q}}[p/p^{*}])\leq 0$, whereas for the reverse implication one divides (4.4) by $t$, which is then sent to zero to yield (4.3). ∎
We continue to identify measures with their $\mu$-densities. Theorem 4.6 can be used to recover the following result of Lardy et al. (2023), who define the maximum description gain of switching from $p\in{\mathcal{P}}$ to some other element of ${\mathcal{P}}$ by
$H({\mathbb{Q}}\mid p\leadsto{\mathcal{P}})=\sup_{p^{\prime}\in{\mathcal{P}}}{\mathbb{E}}_{\mathbb{Q}}\left[\log\frac{p^{\prime}}{p}\right]$
if ${\mathbb{Q}}$ is absolutely continuous with respect to $p$, and +$\infty$+ ∞ otherwise. The following two corollaries confirm, in particular, that the definition of the RIPr used in this paper, Definition 3.2, is consistent with existing definitions in the literature.
@@
If ${\mathcal{P}}$ is convex and $(p_{n})_{n\in{\mathbb{N}}}$ is a sequence in ${\mathcal{P}}$ with $H({\mathbb{Q}}\mid p_{n}\leadsto{\mathcal{P}})\to 0$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathbb{Q}})$, where $p^{*}$ is the RIPr.
@@
The key observation of Lardy et al. (2023) is that the description gain controls the $L^{1}({\mathbb{Q}})$ distance between the log-densities $\log p_{n}$, showing that they form a Cauchy sequence. We give a short self-contained version of their argument. We claim that for any $t>0$ one has
@@
Setting $t=p_{m}/p_{n}$ in (4.5), taking expectation under ${\mathbb{Q}}$, and using that ${\mathcal{P}}$ is convex and hence contains $(p_{m}+p_{n})/2$, we get
$\frac{1}{9}{\mathbb{E}}_{\mathbb{Q}}\left[\left|\log\frac{p_{m}}{p_{n}}\right|\wedge\left|\log\frac{p_{m}}{p_{n}}\right|^{2}\right]\leq{\mathbb{E}}_{\mathbb{Q}}\left[\log\left(\frac{p_{m}+p_{n}}{2p_{m}}\right)\right]+{\mathbb{E}}_{\mathbb{Q}}\left[\log\left(\frac{p_{m}+p_{n}}{2p_{n}}\right)\right]\leq H_{m}+H_{n},$
where we use the shorthand notation $H_{n}=H({\mathbb{Q}}\mid p_{n}\leadsto{\mathcal{P}})$. From the general inequality ${\mathbb{E}}[Y]\leq{\mathbb{E}}[Y\wedge Y^{2}]+\sqrt{{\mathbb{E}}[Y\wedge Y^{2}]}$ for nonnegative random variables $Y$ we then deduce
${\mathbb{E}}_{\mathbb{Q}}\left[|\log p_{m}-\log p_{n}|\right]\leq 9(H_{m}+H_{n})+3\sqrt{H_{m}+H_{n}}.$
This shows that $\log p_{n}$, $n\in{\mathbb{N}}$, is a Cauchy sequence in $L^{1}({\mathbb{Q}})$ and converges to some limit which we may write as $\log p^{*}$ for some ${\mathbb{Q}}$-almost surely positive random variable $p^{*}$. To ensure that $p^{*}$ is the density of a measure that is equivalent to ${\mathbb{Q}}$ we choose it to be zero on the ${\mathbb{Q}}$-nullset $\{q=0\}$, where $q=d{\mathbb{Q}}/d\mu$. This is easily achieved by replacing $p^{*}$ with $p^{*}\bm{1}_{\{q>0\}}$ if necessary. Now, convergence in $L^{1}$ implies convergence in probability, which is preserved under continuous transformations, so $p_{n}$ converges to $p^{*}$ in probability under ${\mathbb{Q}}$. Furthermore, for any $t\in(0,1)$ and $p\in{\mathcal{P}}$ we have $\log\frac{(1-t)p_{n}+tp}{p_{n}}\geq\log(1-t)>-\infty$, so Fatou’s lemma yields
${\mathbb{E}}_{\mathbb{Q}}\left[\log\frac{(1-t)p^{*}+tp}{p^{*}}\right]\leq\liminf_{n\to\infty}{\mathbb{E}}_{\mathbb{Q}}\left[\log\frac{(1-t)p_{n}+tp}{p_{n}}\right]\leq\lim_{n\to\infty}H_{n}=0\text{ for all }p\in{\mathcal{P}}.$
Theorem 4.6 now shows that $p^{*}$ is the RIPr, provided we can argue that $p^{*}$ belongs to ${\mathcal{P}}^{\circ\circ}$. This will use that $p^{*}=0$ on $\{q=0\}$. Specifically, note that ${\mathcal{P}}^{\circ\circ}$ contains $p_{n}$ and hence, being solid, also the smaller random variable $p_{n}\bm{1}_{\{q>0\}}$. Convergence in ${\mathbb{Q}}$-probability of $p_{n}$ to $p^{*}$ implies convergence in $\mu$-probability of $p_{n}\bm{1}_{\{q>0\}}$ to $p^{*}\bm{1}_{\{q>0\}}=p^{*}$. Since ${\mathcal{P}}^{\circ\circ}$ is closed in $\mu$-probability, it must contain $p^{*}$. This completes the proof. ∎
The maximum description gain is useful in situations where the relative entropy $H({\mathbb{Q}}\mid p)$ is infinite for all $p\in{\mathcal{P}}$. However, in many cases the relative entropies are finite and one recovers the original result of Li (1999) on existence of the reverse information projection, as was also observed by Lardy et al. (2023).
@@
If ${\mathcal{P}}$ is convex, $H({\mathbb{Q}}\mid p)<\infty$ for some $p\in{\mathcal{P}}$, and $(p_{n})_{n\in{\mathbb{N}}}$ is an entropy minimizing sequence in ${\mathcal{P}}$, meaning that $H({\mathbb{Q}}\mid p_{n})\to\inf_{p\in{\mathcal{P}}}H({\mathbb{Q}}\mid p)$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathbb{Q}})$, where $p^{*}$ is the RIPr.
@@
Since the quantities involved are finite for $n$ large enough, $\lim_{n\to\infty}H({\mathbb{Q}}\mid p_{n}\leadsto{\mathcal{P}})=\lim_{n\to\infty}H({\mathbb{Q}}\mid p_{n})-\inf_{p\in{\mathcal{P}}}H({\mathbb{Q}}\mid p)=0$, and the result follows from Corollary 4.8. ∎

### 22 Mar 2024 v2

@@ 4 Finding the numeraire
We identify ${\mathcal{P}}$ with its set of densities $C=\{p\in L^{0}_{+}(\mu)\colon p=d{\mathsf{P}}/d\mu,\ {\mathsf{P}}\in{\mathcal{P}}\}$. The polar of $C$ in $L^{0}_{+}(\mu)$ is $C^{\circ}=\{f\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }p\in C\}$, which is the set of (equivalence classes of) random variables that are $\mu$-almost surely equal to a finite e-variable. We claim that all elements of ${\mathcal{P}}^{\circ\circ}$ are absolutely continuous with respect to $\mu$, and that its set of densities is precisely the bipolar $C^{\circ\circ}=\{p\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }f\in C^{\circ}\}$. To check absolute continuity, let ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ and consider any event $A$ such that $\mu(A)=0$. Then $\infty\bm{1}_{A}$ is an e-variable, hence ${\mathbb{E}}_{\mathsf{P}}[\infty\bm{1}_{A}]\leq 1$, and thus ${\mathsf{P}}(A)=0$. This shows absolute continuity and lets us identify ${\mathcal{P}}^{\circ\circ}$ with its set of densities, namely, the set of $p\in L^{0}_{+}(\mu)$ such that ${\mathbb{E}}_{\mu}[pX]\leq 1$ for all finite e-variables $X$. Here we may restrict to finite e-variables $X$ because ${\mathbb{E}}_{\mu}[pX]\leq 1$ if and only if ${\mathbb{E}}_{\mu}[p(X\wedge n)]\leq 1$ for all $n\in{\mathbb{N}}$, and each $X\wedge n$ is a finite e-variable. These are, up to $\mu$-nullsets, precisely the elements of $C^{\circ}$, so we deduce that ${\mathcal{P}}^{\circ\circ}$ is indeed identified with $C^{\circ\circ}$. Finally, the bipolar theorem of Brannath and Schachermayer (1999, Theorem 1.3) states that $C^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains $C$. This completes the proof. ∎
Corollary 4.6.
If ${\mathcal{P}}$ is finite, then ${\mathcal{P}}^{\circ\circ}$ consists of all sub-probabilities that are set-wise dominated by an element of $\operatorname{conv}({\mathcal{P}})$. In particular, ${\mathcal{P}}^{\circ\circ}\cap M_{1}=\operatorname{conv}({\mathcal{P}})$.
Proof.
For finite ${\mathcal{P}}=\{{\mathsf{P}}_{1},\ldots,{\mathsf{P}}_{n}\}$ we may take $\mu=({\mathsf{P}}_{1}+\cdots+{\mathsf{P}}_{n})/n$ as reference measure. In this case $\operatorname{conv}({\mathcal{P}})$ is already $\mu$-closed and the form of ${\mathcal{P}}^{\circ\circ}$ follows from Lemma 4.5. ∎
@@
If a reference measure $\mu$ is given, we identify any ${\mathsf{P}}\in{\mathcal{P}}$ with its density $p=d{\mathsf{P}}/d\mu$. Thanks to Lemma 4.5, we may similarly identify ${\mathcal{P}}^{\circ\circ}$ with its set of densities. For the sake of brevity we abuse notation and write, for example, $p\in{\mathcal{P}}$ even though, strictly speaking, ${\mathcal{P}}$ is a set of measures, not densities. Next, given ${\mathsf{Q}}$ with density $q$, the absolutely continuous part of any $p\in{\mathcal{P}}^{\circ\circ}$ has density $p^{a}=p\bm{1}_{\{q>0\}}$, and this is ${\mathsf{Q}}$-almost surely equal to $p$ itself. In view of these observations, the conditions (ii) and (iii) of Theorem 3.4 characterizing the RIPr state, respectively, that ${\mathbb{E}}_{\mathsf{Q}}[p/p^{*}]\leq 1$ and ${\mathbb{E}}_{\mathsf{Q}}[\log(p/p^{*})]\leq 0$, for all $p\in{\mathcal{P}}^{\circ\circ}$. Here $p^{*}=d{\mathsf{P}}^{*}/d\mu$. The following theorem simplifies these conditions by showing that, essentially, it is enough to check them for elements of ${\mathcal{P}}$ rather than ${\mathcal{P}}^{\circ\circ}$.
@@
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$ and that ${\mathsf{Q}}$ and all elements of ${\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Let $p^{*}\in{\mathcal{P}}^{\circ\circ}$ be equivalent to ${\mathsf{Q}}$. Then each of the conditions (i), (ii), and (iii) of Theorem 3.4 is equivalent to
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{p}{p^{*}}\right]\leq 1\text{ for all $p\in{\mathcal{P}}$,}$ (4.3)
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{(1-t)p^{*}+tp}{p^{*}}\right]\leq 0\text{ for all $t\in(0,1)$ and $p\in{\mathcal{P}}$.}$ (4.4)
@@
The reason for the asymmetry between Theorem 3.4(iii) and (4.4) is that the latter only involves $p$ from ${\mathcal{P}}$, not ${\mathcal{P}}^{\circ\circ}$. Since ${\mathcal{P}}$ may not contain $p^{*}$, nor be convex, we need to explicitly include convex combinations with $p^{*}$ in the numerator on the left-hand side of (4.4). Let us also emphasize that $t=1$ is excluded in (4.4). This sometimes makes the condition easier to check, for instance in the proof of Corollary 4.9 below.
@@
We know that (i), (ii), and (iii) of Theorem 3.4 are equivalent for elements of ${\mathcal{P}}^{\circ\circ}$ that are equivalent to ${\mathsf{Q}}$. To prove the theorem it is therefore enough to show that that Theorem 3.4(ii), (4.3), and (4.4) are equivalent.
We first show that Theorem 3.4(ii) is equivalent to (4.3). The former is clearly a stronger condition because it states that (4.3) holds with the larger set ${\mathcal{P}}^{\circ\circ}$ in place of ${\mathcal{P}}$. For the converse, suppose (4.3) holds. Let ${\mathcal{P}}^{\prime}$ be the set of all sub-probability densities dominated by some element of the convex hull of ${\mathcal{P}}$, and let ${\mathcal{P}}^{\prime\prime}$ be the closure in $\mu$-probability of ${\mathcal{P}}^{\prime}$. It is clear that (4.3) holds for all $p\in{\mathcal{P}}^{\prime}$ and then, by Fatou’s lemma, also for all $p\in{\mathcal{P}}^{\prime\prime}$. We claim that ${\mathcal{P}}^{\prime\prime}$ is a closed convex solid set that contains ${\mathcal{P}}$; it must then also contain ${\mathcal{P}}^{\circ\circ}$ since this is the smallest such set by Lemma 4.5. From this we conclude that Theorem 3.4(ii) holds. To see that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid, note that by definition, every $p^{\prime\prime}\in{\mathcal{P}}^{\prime\prime}$ is the limit in probability of some $p_{n}^{\prime}\in{\mathcal{P}}^{\prime}$, meaning that $p_{n}^{\prime}\leq p_{n}$ for some $p_{n}$ in the convex hull of ${\mathcal{P}}$. From this description convexity follows directly, closedness is obtained from a diagonal argument, and solidity follows by noting that any $\widetilde{p}\leq p^{\prime\prime}$ is the limit of $\widetilde{p}\wedge p_{n}^{\prime}$, which is still dominated by $p_{n}$ and thus belongs to ${\mathcal{P}}^{\prime}$.
The equivalence of (4.3) and (4.4) is argued as the equivalence of (ii) and (iii) of Theorem 3.4, or the equivalence of the numeraire property and log-optimality in Proposition 2.4. Specifically, the forward implication uses Jensen’s inequality and (4.3) to get ${\mathbb{E}}_{\mathsf{Q}}[\log(1-t+tp/p^{*})]\leq\log(1-t+t{\mathbb{E}}_{\mathsf{Q}}[p/p^{*}])\leq 0$, whereas for the reverse implication one divides (4.4) by $t$, which is then sent to zero to yield (4.3). ∎
We continue to identify measures with their $\mu$-densities. Theorem 4.7 can be used to recover the following result of Lardy et al. (2023), who define the maximum description gain of switching from a particular $p\in{\mathcal{P}}$ to some other element of ${\mathcal{P}}$ by
$H({\mathsf{Q}}\mid p\leadsto{\mathcal{P}})=\sup_{p^{\prime}\in{\mathcal{P}}}{\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{p^{\prime}}{p}\right]$
if ${\mathsf{Q}}$ is absolutely continuous with respect to $p$, and +$\infty$+ ∞ otherwise. The following two corollaries confirm, in particular, that the definition of the RIPr used in this paper, Definition 3.2, is consistent with existing definitions in the literature.
@@
If ${\mathcal{P}}$ is convex and $(p_{n})_{n\in{\mathbb{N}}}$ is a sequence in ${\mathcal{P}}$ with $H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})\to 0$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathsf{Q}})$, where $p^{*}$ is the RIPr.
@@
The key observation of Lardy et al. (2023) is that the description gain controls the $L^{1}({\mathsf{Q}})$ distance between the log-densities $\log p_{n}$, showing that they form a Cauchy sequence. We give a self-contained version of their argument. We claim that for any $t>0$ one has
@@
Setting $t=p_{m}/p_{n}$ in (4.5), taking expectation under ${\mathsf{Q}}$, and using that ${\mathcal{P}}$ is convex and hence contains $(p_{m}+p_{n})/2$, we get
$\frac{1}{9}{\mathbb{E}}_{\mathsf{Q}}\left[\left|\log\frac{p_{m}}{p_{n}}\right|\wedge\left|\log\frac{p_{m}}{p_{n}}\right|^{2}\right]\leq{\mathbb{E}}_{\mathsf{Q}}\left[\log\left(\frac{p_{m}+p_{n}}{2p_{m}}\right)\right]+{\mathbb{E}}_{\mathsf{Q}}\left[\log\left(\frac{p_{m}+p_{n}}{2p_{n}}\right)\right]\leq H_{m}+H_{n},$
where we use the shorthand notation $H_{n}=H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})$. From the general inequality ${\mathbb{E}}[Y]\leq{\mathbb{E}}[Y\wedge Y^{2}]+\sqrt{{\mathbb{E}}[Y\wedge Y^{2}]}$ for nonnegative random variables $Y$ we then deduce
${\mathbb{E}}_{\mathsf{Q}}\left[|\log p_{m}-\log p_{n}|\right]\leq 9(H_{m}+H_{n})+3\sqrt{H_{m}+H_{n}}.$
This shows that $\log p_{n}$, $n\in{\mathbb{N}}$, is a Cauchy sequence in $L^{1}({\mathsf{Q}})$ and converges to some limit which we may write as $\log p^{*}$ for some ${\mathsf{Q}}$-almost surely positive random variable $p^{*}$. To ensure that $p^{*}$ is the density of a measure that is equivalent to ${\mathsf{Q}}$ we choose it to be zero on the ${\mathsf{Q}}$-nullset $\{q=0\}$, where $q=d{\mathsf{Q}}/d\mu$. This is easily achieved by replacing $p^{*}$ with $p^{*}\bm{1}_{\{q>0\}}$ if necessary. Now, convergence in $L^{1}$ implies convergence in probability, which is preserved under continuous transformations, so $p_{n}$ converges to $p^{*}$ in probability under ${\mathsf{Q}}$. Furthermore, for any $t\in(0,1)$ and $p\in{\mathcal{P}}$ we have $\log\frac{(1-t)p_{n}+tp}{p_{n}}\geq\log(1-t)>-\infty$, so Fatou’s lemma yields
${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{(1-t)p^{*}+tp}{p^{*}}\right]\leq\liminf_{n\to\infty}{\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{(1-t)p_{n}+tp}{p_{n}}\right]\leq\lim_{n\to\infty}H_{n}=0\text{ for all }p\in{\mathcal{P}}.$
Theorem 4.7 now shows that $p^{*}$ is the RIPr, provided we can argue that $p^{*}$ belongs to ${\mathcal{P}}^{\circ\circ}$. This will use that $p^{*}=0$ on $\{q=0\}$. Specifically, note that ${\mathcal{P}}^{\circ\circ}$ contains $p_{n}$ and hence, being solid, also the smaller random variable $p_{n}\bm{1}_{\{q>0\}}$. Convergence in ${\mathsf{Q}}$-probability of $p_{n}$ to $p^{*}$ implies convergence in $\mu$-probability of $p_{n}\bm{1}_{\{q>0\}}$ to $p^{*}\bm{1}_{\{q>0\}}=p^{*}$. Since ${\mathcal{P}}^{\circ\circ}$ is closed in $\mu$-probability, it must contain $p^{*}$. This completes the proof. ∎
The maximum description gain is useful in situations where the relative entropy $H({\mathsf{Q}}\mid p)$ is infinite for all $p\in{\mathcal{P}}$. However, in many cases the relative entropies are finite and one recovers the original result of Li (1999) on existence of the reverse information projection, as was also observed by Lardy et al. (2023).
@@
If ${\mathcal{P}}$ is convex, $H({\mathsf{Q}}\mid p)<\infty$ for some $p\in{\mathcal{P}}$, and $(p_{n})_{n\in{\mathbb{N}}}$ is an entropy minimizing sequence in ${\mathcal{P}}$, meaning that $H({\mathsf{Q}}\mid p_{n})\to\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathsf{Q}})$, where $p^{*}$ is the RIPr.
@@
Since the quantities involved are finite for $n$ large enough, $\lim_{n\to\infty}H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})=\lim_{n\to\infty}H({\mathsf{Q}}\mid p_{n})-\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)=0$, and the result follows from Corollary 4.9. ∎